OpenStudy (anonymous):

a sum of money is divided among three people in the ratio 15:18:7.find the total sum of money and the largest share,given that the smallest share is $84

OpenStudy (campbell_st):

well the smallest part in the ratio is 7 so then 7 parts = $84 you can use this to find 1 part of the ratio \[1 = \frac{84}{7}\] when you know the value of 1 part you can find the largest share which is 18 parts. so take the value of 1 part and multiply it by 18. hope it helps

OpenStudy (anonymous):

Can you please further explain it........I mean can you write whole of it in a form of equation for me?

OpenStudy (mokeira):

we can change all of them to fractions \[\frac{ 15 }{40 }, \frac{ 18 }{ 40 }, \frac{ 7 }{40 }\] the smallest share looks like it is represented by 7/40 therefore \[\frac{ 7 }{40 }=$84\] to get largest share: 

OpenStudy (mokeira):

to find total sum:

OpenStudy (anonymous):

Thank you very much Mokeira:)

OpenStudy (mokeira):

glad to help